OpenStudy (anonymous):

Find a rational number and an irrational number that are between 7.7 and 7.9. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (anonymous):

@Michele_Laino @Nnesha

OpenStudy (anonymous):

ok so this is weird i dont know how to start or what to do... :(

OpenStudy (michele_laino):

If I call with q the rational number, we have: \[\Large 7.7 = \frac{{77}}{{10}} < q < \frac{{79}}{{10}} = 7.9\]

OpenStudy (anonymous):

can you do this step by step please? i dont understand

OpenStudy (michele_laino):

we can write this: \[7.7 = \frac{{77}}{{10}}\] and \[\frac{{79}}{{10}} = 7.9\]

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (michele_laino):

so a rational number which is between 7.7 and 7.9, also checks this equivalent condition: \[\Large 7.7 < q < 7.9 \Rightarrow \frac{{77}}{{10}} < q < \frac{{79}}{{10}}\]

OpenStudy (anonymous):

oh so wouldnt it have to be 78/10?

OpenStudy (anonymous):

its the only one in between lol

OpenStudy (anonymous):

7.8

OpenStudy (michele_laino):

correct, a possible choice for q is 78/10= 39/5

OpenStudy (anonymous):

ohh ok ill come back later and write the answer but i have another and i think the last question:)

OpenStudy (michele_laino):

we have to find an irrational number